OpenStudy (anonymous):

Choose the correct procedure used to graph the equation 3x + 9y = -18 using the intercepts method.

OpenStudy (abmon98):

Change into slope intercept form. 9y=-3x-18 y=-1/3x-2 solve for both x-intercept and y-intercept. when x=0 y=2 when y=0 2=-x/3 x=-6

OpenStudy (anonymous):

so what do i do ? @Abmon98

OpenStudy (anonymous):

@superhelp101

OpenStudy (abmon98):

read what i have wrote above you will find the x-intercept when y=0 and the y-intercept when x=0